OpenStudy (cutiecomittee123):

The second term in a geometric sequence is 20. The fourth term in the same sequence is 45/4 or 11.25. What is the common ratio in this sequence?

OpenStudy (triciaal):

geometric a ar ar^2 etc Tn = ar^n-1 a = first term ratio = r T4 = 45/4 = ar^3 T2= 20 = ar

OpenStudy (triciaal):